# Overcoming Long-Context Limitations of State-Space Models via Context-Dependent Sparse Attention

### v2

@@ Abstract
Efficient long-context modeling remains a critical challenge for natural language processing (NLP), as the time complexity of the predominant Transformer architecture scales quadratically with the sequence length. While state-space models (SSMs) offer alternative sub-quadratic solutions, they struggle to capture long-range dependencies effectively. In this work, we focus on analyzing and improving the long-context modeling capabilities of SSMs. We show that the widely used synthetic task, associative recall, which requires a model to recall a value associated with a single key without context, insufficiently represents the complexities of real-world long-context modeling. To address this limitation, we extend the associative recall to a novel synthetic task, joint recall, which requires a model to recall the value associated with a key given in a specified context. Theoretically, we prove that SSMs do not have the expressiveness to solve multi-query joint recall in sub-quadratic time complexity. To resolve this issue, we propose a solution based on integrating SSMs with Context-Dependent Sparse Attention (CDSA), which has the expressiveness to solve multi-query joint recall with sub-quadratic computation. To bridge the gap between theoretical analysis and real-world applications, we propose locality-sensitive Hashing Attention with sparse Key Selection (HAX), which instantiates the theoretical solution and is further tailored to natural language domains. Extensive experiments on both synthetic and real-world long-context benchmarks show that HAX consistently outperforms SSM baselines and SSMs integrated with context-independent sparse attention (CISA).
Machine Learning, ICML
@@ 1 Introduction
Long-context modeling is a central challenge in natural language processing (NLP), which underpins a variety of applications, such as document summarization, question answering, and machine translation (Pawar et al., 2024). Recent advances in large language models (LLMs) have broadened the landscape of long-context modeling, enabling new capabilities such as autonomous agents, retrieval-augmented generation, dialogue systems, and long-context reasoning (Liu et al., 2025). This growing demand has spurred intensive research into algorithms that can efficiently and effectively capture long-range dependencies (Tay et al., 2022).
Currently, the Transformer architecture (Vaswani et al., 2017) is the dominant paradigm in sequence modeling. However, its applicability to long sequences is fundamentally constrained by the required computation that grows quadratically with the sequence length. This motivates the research direction for the invention of efficient architectures.
Recently, state-space models (SSMs) (Fu et al., 2023; Gu et al., 2022; Poli et al., 2023) have emerged as a potential alternative solution, offering sub-quadratic time complexity as well as comparable performance to Transformers on short-context NLP tasks (Gu & Dao, 2024; Dao & Gu, 2024). However, empirical evidence suggests that SSMs are less effective than Transformers in capturing long-range dependencies (Waleffe et al., 2024). Furthermore, theoretical analysis by Jelassi et al. (2024) demonstrates that SSMs are much less capable of handling long-context copying, due to the limitations of architecture representation capacity.
In this work, we aim to better understand and improve the long-context modeling abilities of SSMs. We first show that previous studies based on the widely used synthetic task, associative recall (Ba et al., 2016), might be constrained by its limited capability to simulate natural language in-context dependencies. To be specific, associative recall assumes that each key is uniquely associated with a value, regardless of the surrounding context. However, natural language often defies this assumption: the same key can correspond to different values depending on its context. Consider the example in Fig. 1, when asked on which side of the road people drive, the correct answer should depend on the country being referenced. Without specifying whether the context is the US or the UK, the question becomes ambiguous. This example highlights a critical shortcoming of associative recall: it lacks the capacity to simulate context-dependent key-value association, which is very common in natural language.
@@
To overcome this expressiveness bottleneck, we propose to augment SSMs with Context-Dependent Sparse Attention (CDSA), a class of sparse attention with sparse attention patterns that are conditioned on the context representations. Locality-sensitive hashing (LSH) attention (Kitaev et al., 2020) exemplifies CDSA, while context-independent sparse attention (CISA) includes sliding window attention, A-shaped attention, and dilated attention (Ding et al., 2023a). Compared to CISA, CDSA enables dynamic content-dependent routing of information, which is essential to efficiently solve the multi-query joint recall task. We theoretically show that there exists a CDSA which, when integrated with SSMs, enables solving the multi-query joint recall task in sub-quadratic time with respect to sequence length. Moreover, we establish an expressiveness gap between SSMs integrated with CDSA and SSMs integrated with CISA on multi-query joint recall.
Building upon this insight and to bridge the gap between theory and practice, we propose a novel architecture: locality-sensitive Hashing Attention with sparse Key Selection (HAX). HAX improves the expressiveness of LSH attention by incorporating our proposed Key Selection (KS) attention, and is further integrated with state-of-the-art SSMs, Mamba and Mamba2 (Gu & Dao, 2024; Dao & Gu, 2024), instantiating the theoretically grounded solution. We validate the effectiveness of HAX through extensive experiments on both synthetic and real-world long-context modeling benchmarks. The experiment results show that HAX consistently outperforms SSM baselines as well as SSMs augmented with CISA. These findings demonstrate that CDSA, when carefully integrated with SSMs, is a critical component in unlocking their potential for long-context modeling.
@@ 2 Preliminaries
In this section, we introduce two prominent approaches for efficient architecture design: sparse attention in Sec. 2.1 (exemplified by LSH attention in Sec. 2.2) and SSMs in Sec. 2.3. We also introduce associative recall, a widely-used synthetic benchmark for long-context modeling evaluation, in Sec. 2.4.
@@ 2.1 Sparse Attention
with ${\mathbf{S}}\in\{0,1\}^{l\times l}$ representing the sparse attention pattern. To enforce sparsity, we impose the constraint
@@
To ensure per-step computational efficiency, we further tighten this constraint by requiring
@@ 2.2 Locality-Sensitive Hashing Attention
Given that locality-sensitive hashing (LSH) attention represents one of the most effective input-dependent sparse attention for auto-regressive modeling (Kitaev et al., 2020), we reformulate a simple algorithm to generate the sparse attention pattern of LSH. This algorithm accepts the query and key matrices $\mathbf{Q}$} and $\mathbf{K}$} as input and outputs a binary sparse attention pattern ${\mathbf{S}}_{\text{LSH}}$. At each forward pass, $\mathbf{Q}$} and $\mathbf{K}$} are first centralized and normalized to $\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$, respectively:
$\tilde{{\mathbf{Q}}}_{i}=\text{normalize}({\mathbf{Q}}_{i}-\bar{{\mathbf{Q}}}_{i})$ (5)
$\tilde{{\mathbf{K}}}_{i}=\text{normalize}({\mathbf{K}}_{i}-\bar{{\mathbf{K}}}_{i})$ (6)
Next, a random projection matrix ${\mathbf{H}}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)\in\mathbb{R}^{d\times h}$ is sampled to project the normalized vectors $\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$ into the hash space. Then, we consider two binning rules which assign each vector $\tilde{{\mathbf{Q}}}_{i}$ and $\tilde{{\mathbf{K}}}_{i}$ to a discrete hash bin: the argmax binning rule and the sign-bit binning rule.
The argmax binning rule assigns each vector to the index of its most aligned column in $\mathbf{H}$} by dot product:
@@
The sign-bit binning rule constructs a binary hash code by computing the signs of the projected values and interpreting it as a binary number (Ding et al., 2024):
$\text{bin}_{Q_{i}}=\sum_{j=1}^{h}\mathbf{1}[(\tilde{{\mathbf{Q}}}_{i}{\mathbf{H}})_{j}>0]\cdot 2^{h-j}$ (8)
$\text{bin}_{K_{i}}=\sum_{j=1}^{h}\mathbf{1}[(\tilde{{\mathbf{K}}}_{i}{\mathbf{H}})_{j}>0]\cdot 2^{h-j}$ (9)
@@ 2.3 Generalized State-Space Model
Following the definitions introduced by Jelassi et al. (2024), we formulate generalized state-space models as sequence models defined by an update rule $u:\mathcal{U}\times\mathcal{V}\rightarrow\mathcal{U}$ and an output function $r:\mathcal{U}\rightarrow\mathcal{V}$, where $\mathcal{V}$ denotes the token vocabulary and $\mathcal{U}$ represents the recurrent state. Let $U_{0}(\varnothing)\in\mathcal{U}$ denote the initial state. Given an input sequence $v_{1},...,v_{n}\in\mathcal{V}$, for $i$ in $\{1...n\}$, the state $U_{i}(v_{1},...,v_{i})\in\mathcal{U}$ and its corresponding output $R_{i}(v_{1},...,v_{i})\in\mathcal{V}$ are defined recursively as:
@@ 2.4 Associative Recall
The associative recall task was originally introduced in Ba et al. (2016). Olsson et al. (2022) found that the LLM performance on this task is strongly correlated with their in-context learning abilities . Arora et al. (2024) extended associative recall to the multi-query setting: a model is first given a sequence of associated key-value pairs, and then required to recall each value when queried with the corresponding key. Associative recall has been widely adopted as a synthetic benchmark for long-context modeling (Dao & Gu, 2024; Hsieh et al., 2024).
@@ 3.2 Formulation
Associative recall requires a model to memorize $n_{k}$ associated key-value pairs. Joint recall generalizes this task: the model is required to memorize an $n_{c}\times n_{k}$ table of context-specific key-value associations, in which $n_{k}$ keys are associated with different values in each of the $n_{c}$ contexts. Inspired by Arora et al. (2024), we also extend joint recall to a multi-query setting, requiring the model to recover the entire table instead of a specific entry in the table.
Fig. 2 illustrates multi-query joint recall with $n_{c}=2$ and $n_{k}=2$. Following the structure of natural language, the sequentialized table input consists of $n_{c}$ context blocks, each beginning with a context token (e.g. uppercase letter in Fig. 2), followed by $n_{k}$ key-value pairs specific to that context (e.g. lowercase-letter-digit pairs in Fig. 2). Then, the model is tasked with recalling the associated values given each context-key pair, under arbitrary permutations of the context and key ordering.
@@
Figure 2: Comparison between the synthetic formulation of multi-query joint recall and multi-query associative recall.
@@ 3.3.1 Categorization of sparse attention
The sparse attention pattern $\mathbf{S}$} defined in Sec. 2.1 can be categorized as context-dependent or context-independent, depending on whether it is predetermined or dynamically inferred from the context representations. Context-independent sparse attention (CISA) patterns, such as sliding window attention, A-shaped attention, and dilated attention (Ding et al., 2023a), are fixed regardless of context. In contrast, context-dependent sparse attention (CDSA) patterns, exemplified by LSH attention (Kitaev et al., 2020), adapt according to the context representation. Appendix Fig. 9 provides an illustration of both categories.
@@ 3.3.2 Limited expressiveness of SSMs
As a corollary of Theorem 2.7 in Jelassi et al. (2024), we demonstrate that solving the multi-query joint recall task imposes a linear growth requirement on the state dimension of SSMs with respect to the number of entries $n$ in the joint recall table, $n=n_{c}\times n_{k}$. Let $|\mathcal{U}|$ be the number of distinct representations that the recurrent state space $\mathcal{U}$ can encode, for a state with $b$ bits of capacity, $|\mathcal{U}|=2^{b}$. We define the uniform multi-query joint recall distribution as the distribution in which all values are sampled i.i.d. from the uniform distribution over the token vocabulary $\mathcal{V}$. In this setting, we obtain the following:
@@ 4 Method
In this section, we first discuss the expressiveness limitations of LSH attention in Sec. 4.1, and then address these limitations by introducing our proposed key selection (KS) attention in Sec. 4.2. Finally, Sec. 4.3 details the architecture of HAX as well as how HAX is integrated with state-of-the-art SSM architectures, Mamba and Mamba2 (Gu & Dao, 2024; Dao & Gu, 2024).
@@ 4.1 Limitation of LSH Attention
In LLMs, certain keys (particularly those at the beginning of a sequence) often receive attention from most queries, forming distinctive "vertical-stripe" attention patterns (Vig & Belinkov, 2019), as illustrated in Appendix Fig. 10. These globally attended keys play an essential role in instruction following, where the model is expected to focus its attention on the instruction tokens (Liu et al., 2024).
@@
Figure 3: Illustration of the HAX architecture, and Mamba (Gu & Dao, 2024) integrated with HAX.
@@ 5 Experiments
Table 1: Results of Ruler benchmark at 2K context length. We compare different sparse attention integrated with Mamba, including CISA methods: dilated attention (D), sliding window attention (SW), and their combination (SW+D), and A-shaped attention (A), as well as CDSA methods: LSH attention, and our proposed key selection (KS) attention and HAX. The best average performance is in bold.
Mamba 100 95.2 86.8 34.4 2.6 0 32.4 28.35 2.76 1.4 100 25.2 42.43
+D 0 0 0 0 0 0 0 0 17.4 0 100 0 9.78
+SW 100 93.6 83.6 36.2 3.4 0 31.05 31.45 5.24 1.26 100 24.4 42.52
+SW+D 100 95 82.4 34 2.6 0 32.9 29.45 4.72 1.22 100 24.8 42.26
+A 100 93.2 81.4 34.4 2.6 0 31.7 27.1 4.4 1.26 100 26.8 41.91
+LSH 0 0 0 0 0 0 0 0 0.2 0.06 100 0 8.36
+KS (ours) 100 94.4 84.4 33.6 1.8 1.8 37.25 33.85 4.04 1 100 25 43.10
+HAX (ours) 100 98.8 89.6 45.4 3.8 0.2 35.5 28.2 2.04 1.64 100 29.6 44.57
Table 2: Results of LongBench English tasks. We compare different sparse attention integrated with Mamba, including CISA and CDSA methods as in Tab. 1. The best average performance is in bold.
+D 0 0.21 0 2.87 0.26 0 0 0 0 0 0 0.72 3.46 0.7 0 0 0.51
Figure 4: Validation $\mathcal{L}_{LM}$ during continual pre-training. Mamba integrated with HAX is the only architecture that consistently exhibits a decreasing validation loss over the entire training process.

### v3

@@ Abstract
Efficient long-context modeling remains a critical challenge for natural language processing (NLP), as the time complexity of the predominant Transformer architecture scales quadratically with the sequence length. While state-space models (SSMs) offer alternative sub-quadratic solutions, they struggle to capture long-range dependencies effectively. In this work, we focus on analyzing and improving the long-context modeling capabilities of SSMs. We show that the widely used synthetic task, associative recall, which requires a model to recall a value associated with a single key without context, insufficiently represents the complexities of real-world long-context modeling. To address this limitation, we extend the associative recall to a novel synthetic task, joint recall, which requires a model to recall the value associated with a key given in a specified context. Theoretically, we prove that SSMs do not have the expressiveness to solve multi-query joint recall in sub-quadratic time complexity. To resolve this issue, we propose a solution based on integrating SSMs with Context-Dependent Sparse Attention (CDSA), which has the expressiveness to solve multi-query joint recall with sub-quadratic computation. To bridge the gap between theoretical analysis and real-world applications, we propose locality-sensitive Hashing Attention with sparse Key Selection (HAX), which instantiates the theoretical solution and is further tailored to natural language domains. Extensive experiments on both synthetic and real-world long-context benchmarks show that HAX consistently outperforms SSM baselines and SSMs integrated with context-independent sparse attention (CISA). Our code is available at: https://github.com/DeepGraphLearning/HAX.
@@ 1 Introduction
Long-context modeling is a central challenge in natural language processing (NLP), which underpins a variety of applications, such as document summarization, question answering, and machine translation pawar2024whatwhycontextlength . Recent advances in large language models (LLMs) have broadened the landscape of long-context modeling, enabling new capabilities such as autonomous agents, retrieval-augmented generation, dialogue systems, and long-context reasoning liu2025comprehensivesurveylongcontext . This growing demand has spurred intensive research into algorithms that can efficiently and effectively capture long-range dependencies tay2022efficienttransformerssurvey .
Currently, the Transformer architecture transformer is the dominant paradigm in sequence modeling. However, its applicability to long sequences is fundamentally constrained by the required computation that grows quadratically with the sequence length. This motivates the research direction for the invention of efficient architectures.
Recently, state-space models (SSMs) h3 ; s4 ; hyena have emerged as a potential alternative solution, offering sub-quadratic time complexity as well as comparable performance to Transformers on short-context NLP tasks mamba ; mamba2 . However, empirical evidence suggests that SSMs are less effective than Transformers in capturing long-range dependencies waleffe2024empirical . Furthermore, theoretical analysis by jelassi2024repeat demonstrates that SSMs are much less capable of handling long-context copying, due to the limitations of architecture representation capacity.
In this work, we aim to better understand and improve the long-context modeling abilities of SSMs. We first show that previous studies based on the widely used synthetic task, associative recall associative_recall , might be constrained by its limited capability to simulate natural language in-context dependencies. To be specific, associative recall assumes that each key is uniquely associated with a value, regardless of the surrounding context. However, natural language often defies this assumption: the same key can correspond to different values depending on its context. Consider the example in Figure 1, when asked on which side of the road people drive, the correct answer should depend on the country being referenced. Without specifying whether the context is the US or the UK, the question becomes ambiguous. This example highlights a critical shortcoming of associative recall: it lacks the capacity to simulate context-dependent key-value association, which is very common in natural language.
@@
To overcome this expressiveness bottleneck, we propose to augment SSMs with Context-Dependent Sparse Attention (CDSA), a class of sparse attention with sparse attention patterns that are conditioned on the context representations. Locality-sensitive hashing (LSH) attention reformer exemplifies CDSA, while context-independent sparse attention (CISA) includes sliding window attention, A-shaped attention, and dilated attention longnet . Compared to CISA, CDSA enables dynamic content-dependent routing of information, which is essential to efficiently solve the multi-query joint recall task. We theoretically show that there exists a CDSA which, when integrated with SSMs, enables solving the multi-query joint recall task in sub-quadratic time with respect to sequence length. Moreover, we establish an expressiveness gap between SSMs integrated with CDSA and SSMs integrated with CISA on multi-query joint recall.
Building upon this insight and to bridge the gap between theory and practice, we propose a novel architecture: locality-sensitive Hashing Attention with sparse Key Selection (HAX). HAX improves the expressiveness of LSH attention by incorporating our proposed Key Selection (KS) attention, and is further integrated with state-of-the-art SSMs, Mamba and Mamba2 mamba ; mamba2 , instantiating the theoretically grounded solution. We validate the effectiveness of HAX through extensive experiments on both synthetic and real-world long-context modeling benchmarks. The experiment results show that HAX consistently outperforms SSM baselines as well as SSMs augmented with CISA. These findings demonstrate that CDSA, when carefully integrated with SSMs, is a critical component in unlocking their potential for long-context modeling.
@@ 2 Preliminaries
In this section, we introduce two prominent approaches for efficient architecture design: sparse attention in Sec. 2.1 (exemplified by LSH attention in Sec. 2.2) and SSMs in Sec. 2.3. We also introduce associative recall, a widely-used synthetic benchmark for long-context modeling, in Sec. 2.4.
@@ 2.1 Sparse Attention
with ${\mathbf{S}}\in\{0,1\}^{l\times l}$ representing the sparse attention pattern. We consider the sparsity constraint:
@@
To ensure per-step computational efficiency, we further tighten this constraint by requiring:
@@ 2.2 Locality-Sensitive Hashing Attention
Given that locality-sensitive hashing (LSH) attention represents one of the most effective input-dependent sparse attention for auto-regressive modeling reformer , we reformulate a simple algorithm to generate the sparse attention pattern of LSH. This algorithm accepts the query and key matrices $\mathbf{Q}$} and $\mathbf{K}$} as input and outputs a binary sparse attention pattern ${\mathbf{S}}_{\text{LSH}}$. At each forward pass, $\mathbf{Q}$} and $\mathbf{K}$} are first centralized and normalized to $\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$, respectively:
$\tilde{{\mathbf{Q}}}_{i}=\text{normalize}({\mathbf{Q}}_{i}-\bar{{\mathbf{Q}}}_{i}),\quad\tilde{{\mathbf{K}}}_{i}=\text{normalize}({\mathbf{K}}_{i}-\bar{{\mathbf{K}}}_{i})$ (5)
Next, a random projection matrix ${\mathbf{H}}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,1)\in\mathbb{R}^{d\times h}$ is sampled to project the normalized vectors $\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$ into the hash space. Then, we consider two binning rules which assign each vector $\tilde{{\mathbf{Q}}}_{i}$ and $\tilde{{\mathbf{K}}}_{i}$ to a hash bin: the argmax binning rule reformer ; lha and the sign-bit binning rule memoryformer ; hashatt .
The argmax binning rule assigns each vector to the index of its most aligned column in $\mathbf{H}$}:
@@
The sign-bit binning rule constructs a binary hash code by computing the signs of the projected values and interpreting it as a binary number:
$\text{bin}_{Q_{i}}=\sum_{j=1}^{h}\mathbf{1}[(\tilde{{\mathbf{Q}}}_{i}{\mathbf{H}})_{j}>0]\cdot 2^{h-j},\quad\text{bin}_{K_{i}}=\sum_{j=1}^{h}\mathbf{1}[(\tilde{{\mathbf{K}}}_{i}{\mathbf{H}})_{j}>0]\cdot 2^{h-j}$ (7)
@@ 2.3 Generalized State-Space Model
Following the definitions introduced by jelassi2024repeat , we formulate generalized state-space models as sequence models defined by an update rule $u:\mathcal{U}\times\mathcal{V}\rightarrow\mathcal{U}$ and an output function $r:\mathcal{U}\rightarrow\mathcal{V}$, where $\mathcal{V}$ denotes the token vocabulary and $\mathcal{U}$ represents the recurrent state. Let $U_{0}(\varnothing)\in\mathcal{U}$ denote the initial state. Given an input sequence $v_{1},...,v_{n}\in\mathcal{V}$, for $i$ in $\{1...n\}$, the state $U_{i}(v_{1},...,v_{i})\in\mathcal{U}$ and its corresponding output $R_{i}(v_{1},...,v_{i})\in\mathcal{V}$ are defined recursively as:
@@ 2.4 Associative Recall
The associative recall task was originally introduced in associative_recall . olsson2022incontextlearning found that the LLM performance on this task is strongly correlated with their in-context learning abilities . zoology extended associative recall to the multi-query setting: a model is first given a sequence of associated key-value pairs, and then required to recall each value when queried with the corresponding key. Associative recall has been widely adopted as a synthetic benchmark for long-context modeling mamba2 ; ruler .
@@ 3 Joint Recall
Figure 2: Comparison between synthetic multi-query joint recall and associative recall.
@@ 3.2 Formulation
Associative recall requires a model to memorize $n_{k}$ associated key-value pairs. Joint recall generalizes this task: the model is required to memorize an $n_{c}\times n_{k}$ table of context-specific key-value associations, in which $n_{k}$ keys are associated with different values in each of the $n_{c}$ contexts. Inspired by zoology , we also extend joint recall to a multi-query setting, requiring the model to recover the entire table instead of a specific entry in the table.
Figure 2 illustrates multi-query joint recall with $n_{c}=2$ and $n_{k}=2$. Following the structure of natural language, the sequentialized table input consists of $n_{c}$ context blocks, each beginning with a context token (e.g. uppercase letter in Figure 2), followed by $n_{k}$ key-value pairs specific to that context (e.g. lowercase-letter-digit pairs in Figure 2). Then, the model is asked to recall the associated values given each context-key pair, under arbitrary permutations of the context and key ordering.
@@ 3.3.1 Categorization of sparse attention
The sparse attention pattern $\mathbf{S}$} defined in Sec. 2.1 can be categorized as context-dependent or context-independent, depending on whether it is predetermined or dynamically inferred from the context representations. Context-independent sparse attention (CISA) patterns, such as sliding window attention, A-shaped attention, and dilated attention longnet , are fixed regardless of context. In contrast, context-dependent sparse attention (CDSA) patterns, exemplified by LSH attention reformer , adapt according to the context representation. Appendix Figure 9 provides an illustration of both categories.
@@ 3.3.2 Limited expressiveness of SSMs
As a corollary of Theorem 2.7 in jelassi2024repeat , we demonstrate that solving the multi-query joint recall task imposes a linear growth requirement on the state dimension of SSMs with respect to the number of entries $n$ in the joint recall table, $n=n_{c}\times n_{k}$. Let $|\mathcal{U}|$ be the number of distinct representations that the recurrent state space $\mathcal{U}$ can encode, for a state with $b$ bits of capacity, $|\mathcal{U}|=2^{b}$. We define the uniform multi-query joint recall distribution as the distribution in which all values are sampled i.i.d. from the uniform distribution over the token vocabulary $\mathcal{V}$. In this setting, we obtain the following:
@@ 4 Method
In this section, we first discuss the expressiveness limitations of LSH attention in Sec. 4.1, and then address these limitations by introducing our proposed key selection (KS) attention in Sec. 4.2. Finally, Sec. 4.3 details the architecture of HAX as well as how HAX is integrated with state-of-the-art SSM architectures, Mamba and Mamba2 mamba ; mamba2 .
@@ 4.1 Limitation of LSH Attention
In LLMs, certain keys (particularly those at the beginning of a sequence) often receive attention from most queries, forming distinctive "vertical-stripe" attention patterns vig2019analyzing , as illustrated in Appendix Figure 10. These globally attended keys play an essential role in instruction following, where the model is expected to focus its attention on the instruction tokens lost_in_the_middle .
@@
Figure 3: Illustration of the HAX architecture, and Mamba mamba integrated with HAX.
@@ 5 Experiments
Figure 4: Validation $\mathcal{L}_{LM}$ during continual pre-training on Mamba 130M models. Mamba integrated with HAX is the only architecture that consistently exhibits a decreasing validation loss over the entire training process.